OpenStudy (anonymous):

The volume of a regular triangular prism is 90 sqrt of 3 cm^3. If the radius of the inscribed circle in the base is 3 cm, find the altitude of the prism. could you help me to solve that?

OpenStudy (anonymous):

no. the answer should be 3.3333..

OpenStudy (akashdeepdeb):

I am sorry you said TRIANGLUR???? :D hold on then..... :D

OpenStudy (anonymous):

haha its ok :)

OpenStudy (akashdeepdeb):



OpenStudy (akashdeepdeb):

base area = rt 3 * (side)^2/4 rt 3 * 3 = 

OpenStudy (akashdeepdeb):

THIS IS THE UPDATED AND CORRECTED IMAGE!!

OpenStudy (akashdeepdeb):

so base area = area of equilateral triangle!! :D rt 3 * (side)^2/4 rt 3 * 27 = 27 rt 3 So altitude = volume/base area = 90 rt 3/27 rt 3 = 90/27 = 3.333333333333333333333333333333333333333333333333 cm :D Understood the question?

OpenStudy (anonymous):

im sorry ha. what is rt by the way?

OpenStudy (akashdeepdeb):

Under root or square root

OpenStudy (anonymous):

i got it. thank you!

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

i have another problem

OpenStudy (anonymous):

A metal hallow bar whose cross-section and dimensions are shown below weighs 8*10^3 kg/m^3 and measures 2m in length. determine the mass of the metal bar with a square hole section.